### v1

@@ Abstract.
Novel techniques include a bootstrap scheme for the associated random walks despite algebraic obstructions, and a refined treatment of Dani’s correspondence. We also establish non-concentration properties of self-similar measures near algebraic subvarieties of $\mathbb{R}^{d}$.
@@
H.Z. is supported by the startup grant of Soochow University.
@@ 1. Introduction
Let $d\geq 1$ be an integer. Denote by $\operatorname{Sim}(\mathbb{R}^{d})$ the group of similarities of $\mathbb{R}^{d}$, i.e. transformations $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ of the form $\boldsymbol{s}\mapsto\mathtt{r}_{\phi}O_{\phi}\boldsymbol{s}+\mathtt{b}_{\phi}$ where $\mathtt{r}_{\phi}>0$ and $O_{\phi}\in O(d)$ is a linear orthogonal transformation of $\mathbb{R}^{d}$, and $\mathtt{b}_{\phi}\in\mathbb{R}^{d}$. A probability measure $\lambda$ on $\operatorname{Sim}(\mathbb{R}^{d})$ is called a randomized self-similar iterated function system (IFS). We say that $\lambda$ is strongly irreducible if $\mathbb{R}^{d}$ is the only finite union of affine subspaces of $\mathbb{R}^{d}$ which is invariant under $\lambda$-almost every $\phi$. We also say that $\lambda$ has a finite exponential moment if there exists $\varepsilon>0$ such that
@@
Throughout this paper, we consider a probability measure $\sigma$ on $\mathbb{R}^{d}$ which is self-similar in the sense that it is stationary under some randomized self-similar IFS $\lambda$ which is strongly irreducible and has a finite exponential moment. Recall that stationarity means
@@
Although we do not impose the maps in the support of $\lambda$ to be contractive, it turns out that the existence of $\sigma$ implies that $\lambda$ is contractive in average, i.e.
$\int_{\operatorname{Sim}(\mathbb{R}^{d})}\log\mathtt{r}_{\phi}\,\mathrm{d}\lambda(\phi)<0,$
and vice versa, see [12, Theorem 2.5].
@@
On $\mathbb{R}$, classical examples of self-similar measures are the Lebesgue measure, the normalized Hausdorff measure on a missing digit Cantor set, or Bernouilli convolutions. In higher dimension, one may consider powers of missing digit Cantor measures, the normalized Hausdorff measure on a Sierpiński triangle, Sierpiński carpet, etc. In general, every randomized self-similar IFS with finite exponential moment and which is contractive in average, has a unique stationary probability measure [38, Theorem 3.1].
The goal of the paper is to study the Diophantine properties of typical points chosen by a self-similar measure on $\mathbb{R}^{d}$. This topic originates from a question of Mahler [43], asking how well points in the middle-thirds Cantor set can be approximated by rationals. Mahler’s question is later recast by Kleinbock-Lindenstrauss-Weiss [33] who ask whether self-similar measures may satisfy a dichotomy in the spirit of Khintchine’s theorem. We first recall this theorem and present our main result, then we discuss how it connects to earlier works.
Given a non-increasing function $\psi:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$, we write $W(\psi)$ the set of $\psi$-approximable vectors in $\mathbb{R}^{d}$. In other terms, a vector $\boldsymbol{s}\in\mathbb{R}^{d}$ belongs to $W(\psi)$ if for infinitely many $(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\mathbb{N}$, one has
(1.1) $\|q\boldsymbol{s}-\boldsymbol{p}\|_{\infty}<\psi(q),$
@@
Let $\lambda$ be a probability measure on $\operatorname{Sim}(\mathbb{R}^{d})$, and assume $\lambda$ is strongly irreducible with finite exponential moment. Let $\sigma$ be a $\lambda$-stationary probability measure on $\mathbb{R}^{d}$. Let $\psi:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$ be a non-increasing function.
@@
(1.2) $\sigma(W(\psi))=\begin{cases}0&\text{ if }\sum_{q\in\mathbb{N}}\psi(q)^{d}<\infty,\\ 1&\text{ if }\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty.\end{cases}$
@@
Here we use the notation $\llbracket 0,n\rrbracket=[0,n]\cap\mathbb{N}$. An asymptotic estimate of the number of primitive solutions of the Diophantine inequality is also provided, see (8.5).
@@
Let us now explain how the topic has evolved from Mahler’s question in the 80’s to the above ˜1.1. In the early literature, the convergence and divergence aspects of ˜1.1 were addressed separately for specific approximation functions. The first significant result in this direction was obtained by Weiss [56] for $\psi(q)=1/q^{1+\varepsilon}$ and measures on the real line satisfying a certain decay condition, including the middle thirds Cantor measure. This work was later generalized by Kleinbock-Lindenstrauss-Weiss [33] to a broader class of measures on $\mathbb{R}^{d}$, known as friendly measures. Subsequent important developments, adopting similar terminology, include [45, 17, 18]. The study of the case $\psi(q)=\varepsilon/q$ was conducted by Einsiedler-Fishman-Shapira [21] for missing digit Cantor measures, and generalized significantly by Simmons-Weiss [52] to arbitary self-similar measures.
@@
Without restriction on the non-increasing function $\psi$, Khalil-Luethi [28] obtained the Khintchine dichotomy (1.2) under the condition that the self-similar measure $\sigma$ has a large enough dimension, and $\lambda$ is rational, finitely supported, contractive, and satisfies the open set condition. For instance, the Cantor measure with single-missing digit in base $5$ is in this class, but not the middle thirds Cantor measure, see also the subsequent related works [58, 19] which all assume large dimension for $\sigma$ in some sense. In our previous paper [7], we proposed a new approach, based on projection theorems à la Bourgain, which led to the Khintchine dichotomy (1.2) in the case where $d=1$, and with no further restriction, i.e. dealing with any self-similar measure $\sigma$ and any non-increasing function $\psi$. The assumption $d=1$ is however needed at the most crucial dimension bootstrapping step in the proof. We also obtained Schmidt’s counting theorem for primitive solutions under the additional assumption $\psi(q)<1/q$. Pushing one step further, ˜1.1 establishes the Khintchine dichotomy for self-similar measures in arbitrary dimension, along with a full Schmidt counting theorem, i.e without any restriction on $\psi$ nor asking solutions of (1.1) to be primitive.
@@
Other related topics. Mahler [43] also suggested the study of intrinsic Diophantine approximations on fractal sets, i.e. approximation by rationals sitting in the fractal itself. For works in this direction, see e.g. [54, 14]. Beside fractal sets, Diophantine approximation on embedded submanifolds has also attracted much attention over the past years, see e.g. [36, 55, 9, 10].
@@
˜1.1 will be deduced from a dynamical statement which we now present. Let $G=\operatorname{SL}_{d+1}(\mathbb{R})$, let $\Lambda\subseteq G$ be a lattice. Consider the quotient space $X=G/\Lambda$ equipped with the standard Riemannian metric and Haar probability measure $m_{X}$. For $x\in X$, we write $\operatorname{inj}(x)$ the injectivity radius of $x$. For $l\in\mathbb{N}$, we denote by $B^{\infty}_{\infty,l}(X)$ the collection of smooth functions on $X$ whose derivatives up to order $l$ are bounded, and set ${\mathcal{S}}_{\infty,l}(\cdot)$ the associated norm (see Section˜2 for more details on these conventions). For $t>0$ and $\boldsymbol{s}=(s_{1},\cdots,s_{d})\in\mathbb{R}^{d}$, consider $a(t),u(\boldsymbol{s})\in G$ given by
@@
˜1.2 establishes the exponential equidistribution of the measure $\sigma$, viewed along a unipotent orbit based at an arbitrary point $x$ and expanded under the associated diagonal flow. The term $\operatorname{inj}(x)^{-1}$ in the error term reflects that equidistribution takes longer when the basepoint $x$ is high in a cusp.
The qualitative convergence (i.e. without rate) implied by ˜1.2 resonates with Khalil-Luethi-Weiss [29], which establishes equidistribution under expansion by a broader class of diagonal flows, but at the cost of restricting the IFS to be “carpet”, i.e. rational, with equal contraction ratios, and no rotation part.
From a quantitative point of view, ˜1.2 can be seen as an effective version of Ratner’s theorem for the multiparameter unipotent flow $(u(\boldsymbol{s}))_{\boldsymbol{s}\in\mathbb{R}^{d}}$ in the context of fractal measures. The non-fractal case (i.e. $\sigma$ is absolutely continuous with respect to Lebesgue) is due to Kleinbock-Margulis, in the broader context of expanding translates of horospherical subgroups [35, 37]. Related works in this direction for the non-horospherical case include [53, 15, 32, 41, 57, 42]. In the context of fractal measures, Khalil-Luethi [28] obtained ˜1.2 under the additional assumption the point $x$ lies in a certain countable set determined by the IFS. They also required the IFS to be rational, contractive, finitely supported, to satisfy the open set condition and has large enough dimension. See also [19] for a different approach for $d=1$. Provided $d=1$, those constraints were eliminated in our previous work [7]. We now generalize [7] to arbitrary dimensions, achieving the theorem in full generality.
The connection between ˜1.2 and the Khintchine dichotomy in ˜1.1 comes from Dani’s philosophy [16]. Roughly speaking, it claims that the Diophantine properties of a vector $\boldsymbol{s}\in\mathbb{R}^{d}$ can be read in the dynamics of the trajectory $(a(t)u(\boldsymbol{s})x_{0})_{t>1}$ on $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$, where $x_{0}$ is the identity coset. The correspondence is rooted in the simple computation
@@
which yields that for every $I\subseteq\mathbb{R},J\subseteq\mathbb{R}^{d}$, the statement
@@
is equivalent to the lattice $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}\subseteq\mathbb{R}^{d+1}$ intersecting the product set $t^{\frac{1}{d+1}}J\times t^{-\frac{d}{d+1}}I$. Now, identifying $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ to the space of covolume $1$ lattices in $\mathbb{R}^{d+1}$ and provided the product set $t^{\frac{1}{d+1}}J\times t^{-\frac{d}{d+1}}I$ looks like a ball, the latter condition can be interpreted dynamically, at which point ˜1.2 can be used. Since Dani’s insight [16], many works have exploited this connection, e.g. [36, 34, 15, 28, 7]. In Section˜8, we will show how effective decorrelation of expanding translates (consequence of ˜1.2) can be utilized to obtain the divergent case of the Khintchine dichotomy, along with a rate as in Schmidt’s counting theorem. This extends [7] which assumed $d=1$ and restricted counting to primitive solutions, both assumptions being required to deal with bounded Siegel transforms. In this paper, we will tackle Siegel transforms which are not even $L^{2}$.
To prove ˜1.2, we exploit the self-similarity of $\sigma$ to see that the translate $a(t)u(\boldsymbol{s})x\,\mathrm{d}\sigma(s)$ is roughly given by the $\log t$-step of a random walk on $X$ associated to $\lambda$ (see Lemma˜7.2). This change of view point originates in the work of Simmons-Weiss [52], see also [46, 47, 28, 20, 1] for further developments in this direction. The random walk is defined as follows. Assume $\lambda$-a.e. $\phi$ is orientation preserving (i.e. $\det O_{\phi}=1$), write $k_{\phi}=\begin{pmatrix}O_{\phi}&\\ &1\end{pmatrix}$, and let $\mu$ be the probability measure on $\operatorname{SL}_{d+1}(\mathbb{R})$ given by
(1.3) $\mu=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\delta_{k^{-1}_{\phi}a(\mathtt{r}^{-1}_{\phi})u(\mathtt{b}_{\phi})}\,\mathrm{d}\lambda(\phi).$
@@
The first phase is to show that the random walk acquires some (small) positive dimension (Proposition˜5.1): there exists $A,\kappa>0$ depending on $\Lambda,\lambda$ such that for any $\rho>0$ small, $x,y\in X$, and any $n\geq\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}x\rvert$, we have
@@
The proof of this statement closely follows the argument in [7, Proposition 3.1], and is based on effective recurrence of the random walk (Proposition˜4.1).
@@
In the second phase, we bootstrap the initial dimension $\kappa$ arbitrarily close to the dimension of the ambient space by convolving with $\mu$ suitably many times (Proposition˜6.2). However, unlike the one-dimensional case in [7], the multislicing method proposed in [6, Section 2] cannot be applied directly to implement this bootstrapping argument in the current higher dimensional setting. This limitation arises because the non-concentration hypothesis described in [6, Theorem 2.1] is never satisfied for $d\geq 2$ due to algebraic obstructions (see Lemma˜6.3). To resolve this limitation, we promote a mild non-concentration hypothesis (MNC) which also enables the dimensional bootstrapping (Definition˜6.9, Proposition˜6.14)  [^1] Such a strategy resonates with the concurrent and independent work of Zuo Lin [40], which manages algebraic obstructions to dimensional bootstrapping in the context of homogeneous spaces of $\operatorname{SL}_{4}(\mathbb{R})$ acted upon by certain two-parameter unipotent flows. . To validate (MNC) in our setting, we first need to rule out potential algebraic obstructions, which is done in Propositions 6.6, 6.7. We then upgrade the absence of obstruction to the non-concentration property (MNC). That part of the argument requires regularity properties of the self-similar measure $\sigma$, namely that $\sigma$ is Hölder regular with respect to proper algebraic subvarieties (˜3.3). We establish this property in Section˜3; it is of independent interest and generalizes some of the results in [33, Section 7].
@@
In the concluding phase, as the dimension approaches that of the ambient space, we complete the proof by applying the spectral gap property of the convolution operator $f\mapsto\mu*f$ acting on $L^{2}(X)$. Finally, ˜1.2 follows from ˜1.3, via the connection between random walks and expanding translates of self-similar measures given in Lemma˜7.2.
@@
To derive the Khintchine dichotomy and Schmidt’s counting theorem, we remark that the convergence part follows from the single effective equidistribution theorem in ˜1.2, as explained in [28, Theorem 9.1]. In order to handle the asymptotic counting in the divergence part, we need to truncate the associated Siegel’s transform appropriately, and then apply different counting strategies according to the values of $\psi$. More precisely, for the case where $\psi(q)$ is not too large, we adopt a refined version of the counting method in [7], which is based on effective double equidistribution and is inspired by the original papers of Schmidt [49, 50]. To handle the part where $\psi(q)$ is large, inspired by Huang-Saxcé [27] and Pfitscher [44], we make use of the fact that for $\sigma$-a.e. $\boldsymbol{s}$, the lattice $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}$ is not too "distorted" as $t\to+\infty$, which guarantees the number of lattice points in $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}$ intersecting a large ball is asymptotic to the volume of the ball. This analysis is conducted in Section˜8. The main challenge for this section, which is new compared to [7], is that we have to deal with Siegel transforms which are not bounded, and not even $L^{2}$ when $d=1$.
@@
Structure of the paper. In Section˜2, we set up notations for the rest of the paper, and recall some basic facts on self-similar measures. In Section˜3, we prove the Hölder regularity of self-similar measures with respect to proper algebraic subvarieties of $\mathbb{R}^{d}$. In Section˜4, we recall the effective recurrence property of the $\mu$-random walk on $X$. In Section˜5, we show that the distribution $\mu^{*n}*\delta_{x}$ acquires small positive dimension at exponentially small scales as long as $n$ is large enough depending on $\operatorname{inj}(x)$. In Section˜6, we promote a mild non-concentration property and implement it in our context to bootstrap the dimension arbitrarily close to $\dim X$. In Section˜7, we deduce Theorems 1.2, 1.3 using the spectral gap of associated Markov operator, we also derive a double equidistribution property. Finally in Section˜8, we show ˜1.1.
@@
Acknowledgements. The authors thank René Pfitscher and Shucheng Yu for insightful discussions concerning Schmidt’s counting theorem and moment estimates of Siegel transforms.

### v2

@@ Abstract.
Novel techniques include a bootstrap scheme for the associated random walks despite algebraic obstructions, and a refined treatment of Dani’s correspondence. Along the way, we also establish non-concentration properties of self-similar measures near algebraic subvarieties of $\mathbb{R}^{d}$.
@@
H.Z. is supported by the startup grant of Soochow University and the National Natural Science Foundation of China (No. $1250$)
@@ 1. Introduction
Let $d\geq 1$ be an integer. Denote by $\operatorname{Sim}(\mathbb{R}^{d})$ the group of similarities of $\mathbb{R}^{d}$, i.e., transformations $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ of the form $\boldsymbol{s}\mapsto\mathtt{r}_{\phi}O_{\phi}\boldsymbol{s}+\mathtt{b}_{\phi}$ where $\mathtt{r}_{\phi}>0$, $O_{\phi}\in O(d)$ is a linear orthogonal transformation of $\mathbb{R}^{d}$, and $\mathtt{b}_{\phi}\in\mathbb{R}^{d}$. A probability measure $\lambda$ on $\operatorname{Sim}(\mathbb{R}^{d})$ is called a randomized self-similar IFS where IFS stands for iterated function system. We will always assume that $\lambda$ is irreducible, i.e. no proper affine subspace of $\mathbb{R}^{d}$ is invariant under $\lambda$-almost every $\phi$. We also suppose that $\lambda$ has a finite exponential moment, meaning there exists $\varepsilon>0$ such that
@@
A probability measure $\sigma$ on $\mathbb{R}^{d}$ is called $\lambda$-stationary if it satisfies
@@
By [12, Theorem 2.5], there exists such a measure $\sigma$ if [^1] In fact the “if” part does not require irreducibility, see [46, Theorem 3.1]. Moreover, whether $\lambda$ is irreducible or not can then be read in the associated stationary measure $\sigma$, as it amounts to $\operatorname{supp}\sigma$ not being included in any affine hyperplane (the direct implication is by Lemma 4.1, the converse is trivial). and only if $\lambda$ is contractive in average, i.e.
$\int_{\operatorname{Sim}(\mathbb{R}^{d})}\log\mathtt{r}_{\phi}\,\mathrm{d}\lambda(\phi)<0.$
In this case, the $\lambda$-stationary probability measure on $\mathbb{R}^{d}$ is unique [12]. Throughout this paper, we consider a probability measure $\sigma$ on $\mathbb{R}^{d}$ which is self-similar in the sense that it is stationary under some randomized self-similar IFS $\lambda$ which is irreducible and has a finite exponential moment.
@@
On the real line $\mathbb{R}$, classical examples of self-similar measures include the Lebesgue measure, the normalized Hausdorff measure on a missing digit Cantor set, or Bernouilli convolutions. In higher dimension, one may consider powers of missing digit Cantor measures, the normalized Hausdorff measure on a Sierpiński triangle, Sierpiński carpet, etc.
The goal of the paper is to study the Diophantine properties of typical points distributed according to a self-similar measure on $\mathbb{R}^{d}$. This topic originates from a question of Mahler [50], asking how well points in the middle-thirds Cantor set can be approximated by rationals. Mahler’s question is later recast by Kleinbock-Lindenstrauss-Weiss [43] who ask whether self-similar measures may satisfy a dichotomy in the spirit of Khintchine’s theorem. Building upon earlier works, we provide an affirmative answer.
Let us first recall the classical Khintchine theorem. Given a non-increasing function $\psi:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$, let us write $W(\psi)$ the set of $\psi$-approximable vectors in $\mathbb{R}^{d}$. In other terms, a vector $\boldsymbol{s}\in\mathbb{R}^{d}$ belongs to $W(\psi)$ if for infinitely many $(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\mathbb{N}$, one has
$\|q\boldsymbol{s}-\boldsymbol{p}\|_{\infty}<\psi(q),$ (1.1)
@@
Let $\lambda$ be a probability measure on $\operatorname{Sim}(\mathbb{R}^{d})$, and assume $\lambda$ is irreducible with finite exponential moment. Let $\sigma$ be a $\lambda$-stationary probability measure on $\mathbb{R}^{d}$. Let $\psi:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$ be a non-increasing function.
@@
$\sigma(W(\psi))=\begin{cases}0&\text{ if }\sum_{q\in\mathbb{N}}\psi(q)^{d}<\infty,\\ 1&\text{ if }\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty.\end{cases}$ (1.2)
@@
Here we use the notation $\llbracket 0,n\rrbracket=[0,n]\cap\mathbb{N}$, and given a finite set $E$, we write $|E|$ for its cardinality. An asymptotic estimate of the number of primitive solutions of the Diophantine inequality is also provided, see (9.5).
@@
Let us now explain how the topic has evolved since Mahler’s question in the 80’s to the above ˜1.1. In the early literature, the convergence and divergence aspects of ˜1.1 were addressed separately for specific approximation functions. The first significant result in this direction was obtained by Weiss [64] for $\psi(q)=1/q^{1+\varepsilon}$ and measures on the real line satisfying a certain decay condition, including the middle thirds Cantor measure. This work was later generalized by Kleinbock-Lindenstrauss-Weiss [43] to a broader class of measures on $\mathbb{R}^{d}$, known as friendly measures. Subsequent important developments, adopting similar terminology, include [54, 18, 19, 16]. The study of the case $\psi(q)=\varepsilon/q$ was conducted by Einsiedler-Fishman-Shapira [24] for missing digit Cantor measures, and generalized significantly by Simmons-Weiss [60] to arbitary self-similar measures.
@@
Without restriction on the non-increasing function $\psi$, Khalil-Luethi [36] obtained the Khintchine dichotomy (1.2) under the condition that the self-similar measure $\sigma$ has a large enough dimension, and $\lambda$ is rational, finitely supported, contractive, and satisfies the open set condition. For instance, the Cantor measure with single-missing digit in base $5$ is in this class, but not the middle thirds Cantor measure, see also the related works [66, 20], which all assume large dimension for $\sigma$ in some sense. In our previous paper [3], we proposed a new approach, based on projection theorems à la Bourgain, which led to the Khintchine dichotomy (1.2) in the case where $d=1$, and with no further restriction, i.e., dealing with any self-similar measure $\sigma$ and any non-increasing function $\psi$. The assumption $d=1$ was however needed at the most crucial step of dimension bootstrapping in the proof. We also obtained Schmidt’s counting theorem for primitive solutions under the additional assumption $\psi(q)<1/q$. Pushing one step further, ˜1.1 establishes the Khintchine dichotomy for self-similar measures in arbitrary dimension, along with a full Schmidt counting theorem, i.e., without any restriction on $\psi$ nor asking solutions of (1.1) to be primitive.
@@
Other related topics. Mahler [50] also suggested the study of intrinsic Diophantine approximations on fractal sets, i.e., approximation by rationals sitting in the fractal itself. For works in this direction, see e.g. [62, 14]. Beside fractal sets, Diophantine approximation on embedded submanifolds has also attracted much attention over the past years, see e.g. [41, 63, 10, 9]. See also [22, 53, 31] for a Khintchine’s theorem and other results of Diophantine approximation on the Grassmannians and other flag varieties.
@@
˜1.1 will be deduced from a dynamical statement which we now present. Let $G=\operatorname{SL}_{d+1}(\mathbb{R})$ and let $\Lambda\subseteq G$ be a lattice. Consider the quotient space $X=G/\Lambda$ and equip it with the Haar probability measure $m_{X}$ and a canonical Riemannian metric (more precisely defined in Section˜2). For $x\in X$, we denote by $\operatorname{inj}(x)$ the injectivity radius of $x$. For $l\in\mathbb{N}$, let $B^{\infty}_{\infty,l}(X)$ be the collection of smooth functions on $X$ all of whose derivatives up to order $l$ are bounded, and set ${\mathcal{S}}_{\infty,l}(\cdot)$ the associated $C^{l}$-norm (see Section˜2 for more details on these conventions). For $t>0$ and $\boldsymbol{s}=(s_{1},\cdots,s_{d})\in\mathbb{R}^{d}$, consider $a(t),u(\boldsymbol{s})\in G$ given by
@@
˜1.2 establishes the exponential equidistribution of the measure $\sigma$, viewed along a unipotent orbit based at an arbitrary point $x$ and expanded under the associated expanding diagonal flow. The term $\operatorname{inj}(x)^{-1}$ in the error term reflects that equidistribution takes longer when the basepoint $x$ is high in a cusp.
The qualitative convergence (i.e. without rate) implied by ˜1.2 resonates with Khalil-Luethi-Weiss [35], which establishes qualitative equidistribution under expansion by a broader class of diagonal flows, but at the cost of restricting the IFS to be “carpet”, i.e. rational with equal contraction ratios and no rotation part.
From a quantitative point of view, ˜1.2 can be seen as an effective version of Ratner’s theorem for the multiparameter unipotent flow $(u(\boldsymbol{s}))_{\boldsymbol{s}\in\mathbb{R}^{d}}$ in the context of fractal measures. The non-fractal case (i.e. $\sigma$ is absolutely continuous with respect to Lebesgue) is due to Kleinbock-Margulis, in the broader context of expanding translates of horospherical subgroups [40, 42], see also [44] for the explicit error term involving injectivity radius. Related works in this direction for the non-horospherical case include [61, 15, 39, 49, 65, 48]. In the context of fractal measures, Khalil-Luethi [36] obtained ˜1.2 under the additional assumption the point $x$ lies in a certain countable set determined by the IFS. They also required the IFS to be rational, contractive, finitely supported, to satisfy the open set condition and has large enough dimension. See also [20] for a different approach for $d=1$. Provided $d=1$, those constraints were eliminated in our previous work [3]. We now generalize [3] to arbitrary dimensions, achieving the theorem in full generality.
The connection between ˜1.2 and the Khintchine dichotomy in ˜1.1 comes from Dani’s philosophy [17]: roughly speaking, the Diophantine properties of a vector $\boldsymbol{s}\in\mathbb{R}^{d}$ can be read in the dynamics of the trajectory $(a(t)u(\boldsymbol{s})x_{0})_{t>1}$ on $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$, where $x_{0}$ denotes the identity coset. The correspondence is rooted in the simple computation
@@
which yields that for every $t>0$, $\boldsymbol{s}\in\mathbb{R}^{d}$, $I\subseteq\mathbb{R},J\subseteq\mathbb{R}^{d}$, the statement
@@
is equivalent to the lattice $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}\subseteq\mathbb{R}^{d+1}$ intersecting the product set $t^{\frac{1}{d+1}}J\times t^{-\frac{d}{d+1}}I$. Now, identifying $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ to the space of covolume $1$ lattices in $\mathbb{R}^{d+1}$ and provided the product set $t^{\frac{1}{d+1}}J\times t^{-\frac{d}{d+1}}I$ looks like a ball, the latter condition can be interpreted dynamically, at which point ˜1.2 can be used. Since Dani’s insight [17], many works have exploited this connection, e.g. [41, 45, 15, 36, 3]. In Section˜9, we will show how effective decorrelation of expanding translates (consequence of ˜1.2) can be utilized to obtain the divergent case of the Khintchine dichotomy, along with a rate as in Schmidt’s counting theorem. This extends [3] which assumed $d=1$ and restricted counting to primitive solutions, both assumptions being required to deal with bounded Siegel transforms. In this paper, we will tackle Siegel transforms which are not even $L^{2}$.
To prove ˜1.2, we exploit the self-similarity of $\sigma$ to see that the translate $a(t)u(\boldsymbol{s})x\,\mathrm{d}\sigma(s)$ is roughly given by the $\log t$-step of a random walk on $X$ associated to $\lambda$ (see Lemma˜8.2). This change of view point originates in the work of Simmons-Weiss [60], see also [56, 55, 36, 21, 1] for further developments in this direction. The random walk is defined as follows. Assume $\lambda$-a.e. $\phi$ is orientation preserving (i.e. $\det O_{\phi}=1$), write $k_{\phi}=\begin{pmatrix}O_{\phi}&\\ &1\end{pmatrix}$, and let $\mu$ be the probability measure on $\operatorname{SL}_{d+1}(\mathbb{R})$ given by
$\mu=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\delta_{k^{-1}_{\phi}a(\mathtt{r}^{-1}_{\phi})u(\mathtt{b}_{\phi})}\,\mathrm{d}\lambda(\phi).$ (1.3)
@@
The first phase is to show that the random walk acquires some (small) positive dimension (Proposition˜6.1): there exists $A,\kappa>0$ depending on $\Lambda,\lambda$ such that for any $\rho>0$ small, $x,y\in X$, and any $n\geq\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}x\rvert$, we have
@@
The proof of this statement closely follows the argument in [3, Proposition 3.1], and is based on effective recurrence of the random walk (Proposition˜5.1).
@@
In the second phase, we bootstrap the initial dimension $\kappa$ arbitrarily close to the dimension of the ambient space by convolving with $\mu$ suitably many times (Proposition˜7.2). However, unlike the one-dimensional case in [3], the multislicing method proposed in [4, Section 2] cannot be applied directly to implement this bootstrapping argument in the current higher dimensional setting. This limitation arises because the non-concentration hypothesis described in [4, Theorem 2.1] is never satisfied for $d\geq 2$ due to algebraic obstructions (see Lemma˜7.3). To resolve this limitation, we promote a mild non-concentration hypothesis (MNC) which also enables the dimensional bootstrapping (Definition˜7.9, Proposition˜7.14)  [^2] Such a strategy resonates with the concurrent and independent work of Zuo Lin [47], which manages algebraic obstructions to dimensional bootstrapping in the context of homogeneous spaces of $\operatorname{SL}_{4}(\mathbb{R})$ acted upon by certain two-parameter unipotent flows..
To validate (MNC) in our setting, we first need to rule out potential algebraic obstructions, which is done in Propositions 7.6, 7.7. We then upgrade the absence of obstruction to the non-concentration property (MNC). That part of the argument requires regularity properties of the self-similar measure $\sigma$, namely that $\sigma$ is Hölder regular with respect to proper algebraic subvarieties (˜4.4). We establish this property in Section˜4. It is of independent interest and complements [51, 33, 52, 43, 27, 30].
@@
In the concluding phase, as the dimension approaches that of the ambient space, we complete the proof by applying the spectral gap property of the convolution operator $f\mapsto\mu*f$ acting on $L^{2}(X)$. Finally, ˜1.2 follows from ˜1.3, via the connection between random walks and expanding translates of self-similar measures given in Lemma˜8.2.
@@
To derive the Khintchine dichotomy and Schmidt’s counting theorem, we remark that the convergence part follows from the single effective equidistribution theorem in ˜1.2, as explained in [36, Theorem 9.1]. In order to handle the asymptotic counting in the divergence part, we need to truncate the associated Siegel’s transform appropriately, and then apply different counting strategies according to the values of $\psi$. More precisely, for the case where $\psi(q)$ is not too large, we adopt a refined version of the counting method in [3], which is based on effective double equidistribution and is inspired by the original papers of Schmidt [57, 58]. To handle the part where $\psi(q)$ is large, inspired by Huang-Saxcé [31] and Pfitscher [53], we make use of the fact that for $\sigma$-a.e. $\boldsymbol{s}$, the lattice $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}$ is not too “distorted” as $t\to+\infty$, which guarantees the number of lattice points in $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}$ intersecting a large ball is asymptotic to the volume of the ball. This analysis is conducted in Section˜9. The main challenge for this section, which is new compared to [3], is that we have to deal with Siegel transforms which are not bounded, and not even $L^{2}$ when $d=1$.
@@
Structure of the paper. In Section˜2, we set up notations for the rest of the paper, and recall some basic facts on self-similar measures. In Section˜3 we discuss moment estimates for $\sigma$ or $\lambda$-trajectories, and derive exponential convergence of $(\lambda^{*n}*\delta_{0})_{n\geq 0}$ to $\sigma$. In Section˜4, we prove the Hölder regularity of self-similar measures with respect to proper algebraic subvarieties of $\mathbb{R}^{d}$. In Section˜5, we recall the effective recurrence property of the $\mu$-random walk on $X$. In Section˜6, we show that the distribution $\mu^{*n}*\delta_{x}$ acquires small positive dimension at exponentially small scales as long as $n$ is large enough depending on $\operatorname{inj}(x)$. In Section˜7, we promote a mild non-concentration property and implement it in our context to bootstrap the dimension arbitrarily close to $\dim X$. In Section˜8, we deduce Theorems 1.2, 1.3 using the spectral gap of associated Markov operator, we also derive a double equidistribution property. Finally in Section˜9, we show ˜1.1.
@@
Acknowledgements. We thank René Pfitscher and Shucheng Yu for insightful discussions concerning Schmidt’s counting theorem and moment estimates of Siegel transforms. We also thank Bing Li for informing us about existing literatures around the Hölder estimates of self-similar measures near submanifolds.